OpenStudy (anonymous):

Prove identity: sin(x – y)sin(x + y) = sin^2 x – sin^2 y

OpenStudy (anonymous):

Just expand each of the \( \sin(x – y) \) and \( \sin(x + y) \) and find the product.

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

Glad to help :) Another interesting identity: \[ \cos(x – y)\cos(x + y) = \cos^2 x – sin^2 y \]